A cattle train left Miami and traveled toward New York. 14 hours later a diesel train left traveling at 45 km/h in effort to catch up to the cattle train. After traveling for four hours the diesel train finally caught up. What was the cattle train's average speed?
Answer by checkley71(8403)

45*4=(14+4)X
180=18X
X=180/18
X=10 KM/H IS THE SPEED OF THE CATTLE TRAIN.